# An Overview on Over-the-air Electromagnetic Signal Processing

### v2

@@ III A General Framework for EM Signal Processing
This section presents a unified conceptual framework for ESP, treating the EM processing structure as a feedback space-variant linear system and relating it to modal analysis. This approach creates a connection between EM theory and SP, outlining the theoretical foundations and performance limits concerning potential processing flexibility.
III-A Preliminaries: The Green Operator and the Wavenumber Representation
@@
We consider a generic space $\mathcal{S}_{\text{T}}$ characterized by the monochromatic current density ${\mathsf{J}}({\bf{r}})$ in A/m2, with ${\bf{r}}=r_{x}{\hat{\bf{x}}}+r_{y}{\hat{\bf{y}}}+r_{z}{\hat{\bf{z}}}\in\mathcal{S}_{\text{T}}$ being the position vector and ${\hat{\bf{x}}}$, ${\hat{\bf{y}}}$, and ${\hat{\bf{z}}}$ the 3D unit vectors. The time dependence $e^{\jmath\omega t}$, with $\omega$ being the angular frequency, has been omitted. In general, space $\mathcal{S}_{\text{T}}$ is a volume, but it may also represent a surface or a wire source. The current density ${\mathsf{J}}({\bf{r}})$ can be an impressed current in the case of an active source, such as a transmitting antenna, or an induced/equivalent current in the case of a scattering device. The electrical field ${\mathsf{E}}({\bf{r}})$ [V/m] generated by ${\mathsf{J}}({\bf{r}})$ can be computed under the free-space condition by considering the Helmholtz equation, whose solution is given by [35]
@@
${\mathsf{E}}({\bf{r}})=\int_{\mathcal{S}_{\text{T}}}\mathsf{G}\left({\bf{r}}-{\bf{s}}\right)\,{\mathsf{J}}({\bf{s}})\,\mathrm{d}{\bf{s}}=\left(\mathbb{G}\,{\mathsf{J}}\right)({\bf{r}})\,.$ (1)
@@
Equation (1) implicitly defines the Green’s tensor operator $\mathbb{G}$, with $\mathsf{G}\left({\bf{r}}\right)$ being the dyadic Green function defined as [35]
@@
$\mathsf{G}\left({\bf{r}}\right)=-\jmath\frac{\eta\,e^{-\jmath k_{0}\,|{\bf{r}}|}}{2\lambda|{\bf{r}}|}\cdot\left[\left({\bf{I}}-\hat{{\bf{r}}}\,\hat{{\bf{r}}}^{\mathrm{T}}\right)+\frac{\jmath\lambda}{2\pi|{\bf{r}}|}\left({\bf{I}}-3\hat{{\bf{r}}}\,\hat{{\bf{r}}}^{\mathrm{T}}\right)-\frac{\lambda^{2}}{(2\pi|{\bf{r}}|)^{2}}\left({\bf{I}}-3\hat{{\bf{r}}}\,\hat{{\bf{r}}}^{\mathrm{T}}\right)\right]$ (2)
@@
where $k_{0}=2\pi/\lambda$ is the wavenumber, $\lambda$ is the wavelength, $\hat{{\bf{r}}}={\bf{r}}/|{\bf{r}}|$, and $\eta$ is the free-space impedance. $\mathsf{G}\left({\bf{r}}\right)$ is the solution of the Helmholtz equation for an impulsive current density ${\mathsf{J}}({\bf{r}})={\mathsf{I}}\cdot\delta({\bf{r}})$, with ${\mathsf{I}}$ being the unit dyadic, [^1] The unit dyadic, also known as the identity dyadic or unit tensor, is a mathematical construct in vector and tensor analysis. It is the second-order tensor equivalent of the scalar number 1 in scalar algebra or the identity matrix in linear algebra. and $\delta({\bf{r}})$ the delta Dirac pseudo-function. When $|{\bf{r}}|\gg\lambda$, only the first term in (2) contributes to the EM radiation. The electrical field in the current-free portion of the space must satisfy the homogeneous wave equation [35]
$\nabla^{2}{\mathsf{E}}({\bf{r}})+k_{0}^{2}\,{\mathsf{E}}({\bf{r}})=0$ (3)
@@
For what follows, it is convenient to introduce the representation of the fields in the wavenumber domain ${\bm{\kappa}}=\kappa_{x}{\hat{\bf{x}}}+\kappa_{y}{\hat{\bf{y}}}+\kappa_{z}{\hat{\bf{z}}}$ through the 3D Fourier transform. Specifically, given a generic field $\mathsf{A}({\bf{r}})$, we define
$\widetilde{\mathsf{A}}({\bm{\kappa}})$ $=\mathcal{F}_{3}\left[\mathsf{A}({\bf{r}})\right]=\int_{{\mathbb{R}}^{3}}\mathsf{A}({\bf{r}})\,e^{-\jmath\,{\bm{\kappa}}\bm{\cdot}{\bf{r}}}\,\mathrm{d}{\bf{r}}$ (4)
with $\bm{\cdot}$ denoting the scalar product.
@@
$\left(\kappa_{x}^{2}+\kappa_{y}^{2}+\kappa_{z}^{2}-k_{0}^{2}\right)\widetilde{\mathsf{E}}({\bm{\kappa}})=0$ (5)
@@
where we have considered that $\mathcal{F}_{3}\left[\nabla^{2}\mathsf{A}({\bf{r}})\right]=-|{\bm{\kappa}}|^{2}\,\widetilde{\mathsf{A}}({\bm{\kappa}})$. The structure of (5) implies that $\widetilde{\mathsf{E}}({\bm{\kappa}})=\mathcal{F}_{3}\left[{\mathsf{E}}({\bf{r}})\right]$ must vanish everywhere except on the wavenumber support
=${\cal{E}}=\left\{(\kappa_{x},\kappa_{y},\kappa_{z})\in\mathbb{R}^{3}:\kappa_{x}^{2}+\kappa_{y}^{2}+\kappa_{z}^{2}=k_{0}^{2}\right\}\,.$ (6)
In other words, the EM propagation phenomenon acts as a filter in the wavenumber domain with support in ${\cal{E}}$. Since $\widetilde{\mathsf{E}}({\bm{\kappa}})$ is defined over a support with zero measure in the waveform domain, it is not an ordinary function but takes the form [36]
$\widetilde{\mathsf{E}}({\bm{\kappa}})=\widetilde{\mathsf{E}}_{0}({\bm{\kappa}})\,\delta\left(\kappa_{x}^{2}+\kappa_{y}^{2}+\kappa_{z}^{2}-k_{0}^{2}\right)$ (7)
@@
where $\widetilde{\mathsf{E}}_{0}({\bm{\kappa}})$ defines the angular dependence (direction) of the wave. When applied to the EM field ${\mathsf{E}}({\bf{r}})$ in the presence of current sources, the inverse Fourier representation expresses it in terms of mathematical plane waves [37, 15]
@@
${\mathsf{E}}({\bf{r}})$ $=\mathcal{F}^{-1}_{3}\left[\widetilde{\mathsf{E}}({\bm{\kappa}})\right]=\frac{1}{(2\pi)^{3}}\int_{{\mathbb{R}}^{3}}\widetilde{\mathsf{E}}({\bm{\kappa}})\,e^{\jmath\,{\bm{\kappa}}\bm{\cdot}{\bf{r}}}\,\text{d}{\bm{\kappa}}=\frac{1}{(2\pi)^{3}}\int_{{\mathbb{R}}^{3}}\widetilde{\mathsf{G}}({\bm{\kappa}})\,{\widetilde{\mathsf{J}}}({\bm{\kappa}})\,e^{\jmath\,{\bm{\kappa}}\bm{\cdot}{\bf{r}}}\,\text{d}{\bm{\kappa}}$
@@
$=\frac{\jmath\eta}{k_{0}(2\pi)^{3}}\int_{{\mathbb{R}}^{3}}\frac{{{\bm{\kappa}}}\,{\scriptstyle\mathsf{x}}\,{{{{\bm{\kappa}}}\,{\scriptstyle\mathsf{x}}\,{{\widetilde{\mathsf{J}}}({\bm{\kappa}})}}}}{|{\bm{\kappa}}|^{2}-k_{0}^{2}}\,e^{\jmath\,{\bm{\kappa}}\bm{\cdot}{\bf{r}}}\,\text{d}{\bm{\kappa}}$ (8)
@@
where $\widetilde{\mathsf{G}}({\bm{\kappa}})=\mathcal{F}_{3}\left[\mathsf{G}\left({\bf{r}})\right)\right]$, ${\widetilde{\mathsf{J}}}({\bm{\kappa}})=\mathcal{F}_{3}\left[{\mathsf{J}}({\bf{r}})\right]$, and ${}\,{\scriptstyle\mathsf{x}}\,{}$ denotes the cross product. In fact, for a fixed wavenumber ${\bm{\kappa}}$, $\widetilde{\mathsf{E}}({\bm{\kappa}})\,e^{\jmath\,{\bm{\kappa}}\bm{\cdot}{\bf{r}}}$ in (8) represents a plane wave with “direction” ${\bm{\kappa}}/k_{0}$. It is interesting to emphasize that the generic plane wave $\widetilde{\mathsf{E}}({\bm{\kappa}})\,e^{\jmath\,{\bm{\kappa}}\bm{\cdot}{\bf{r}}}=\widetilde{\mathsf{E}}({\bm{\kappa}})\,e^{\jmath\,\kappa_{x}\,r_{x}}\,e^{\jmath\,\kappa_{y}\,r_{y}}\,e^{\jmath\,\kappa_{z}\,r_{z}}$ in (8) radiates $\forall{\bf{r}}$, i.e., in all the 3D space, only when $\kappa_{x}$, $\kappa_{y}$, and $\kappa_{z}$ are all real (visible region). Vice versa, if at least one of them is complex, [^2] For example, this might happen if $\kappa_{x}^{2}+\kappa_{y}^{2}>k_{0}^{2}$ so that to satisfy (6) it should be $\kappa_{z}=\jmath\sqrt{\kappa_{x}^{2}+\kappa_{y}^{2}-k_{0}^{2}}\in\mathbb{C}$. let us say $\kappa_{z}$, it is $e^{\jmath\,\kappa_{z}\,r_{z}}=e^{-|\kappa_{z}\,r_{z}|}\rightarrow 0$ for $|r_{z}|$ larger than a few wavelengths and the wave becomes evanescent, i.e., it does not radiate (invisible region). In summary, only the spectral components of $\widetilde{\mathsf{E}}({\bm{\kappa}})$ for ${\bm{\kappa}}\in{\cal{E}}$ represent EM waves, i.e., are physically consistent, and only the subset having real components radiates. For each wavenumber, i.e., plane wave, because of the ${{\bm{\kappa}}}\,{\scriptstyle\mathsf{x}}\,{{{{\bm{\kappa}}}\,{\scriptstyle\mathsf{x}}\,{{\widetilde{\mathsf{J}}}({\bm{\kappa}})}}}$ term in (8), $\widetilde{\mathsf{E}}({\bm{\kappa}})$ is subject to the relation $\widetilde{\mathsf{E}}({\bm{\kappa}})\bm{\cdot}{\bm{\kappa}}=0$, limiting to two the degrees-of-freedom (DoF) achievable with different polarizations [37]. Incidentally, for locations ${\bf{r}}$ in the far-field region of the source, (8) can be approximated as [35, Ch. 12]
${\mathsf{E}}({\bf{r}})$ $\simeq\jmath k_{0}\eta\frac{e^{-\jmath\,k_{0}\,|{\bf{r}}|}}{4\pi|{\bf{r}}|}\,{\hat{\bf{r}}}\,{\scriptstyle\mathsf{x}}\,{\,{{\hat{\bf{r}}}\,{\scriptstyle\mathsf{x}}\,{{\widetilde{\mathsf{J}}}(k_{0}\,\hat{\bf{r}})}}}$ (9)
@@
so that the EM field becomes directly related to the Fourier transform of the source evaluated at direction $k_{0}\,\hat{\bf{r}}\in{\cal{E}}$.

### v3

@@ III A General Framework for EM Signal Processing
III-A Preliminaries: The Green’s Operator and the Wavenumber Representation
We consider a generic source space $\mathcal{S}_{\text{T}}$, characterized by a monochromatic current density ${\mathsf{J}}({\mathbf{r}})$ $\mathrm{A/m^{2}}$], where ${\mathbf{r}}=r_{x}{\hat{\mathbf{x}}}+r_{y}{\hat{\mathbf{y}}}+r_{z}{\hat{\mathbf{z}}}\in\mathcal{S}_{\text{T}}$ denotes the position vector, and $^{\hat{\mathbf{x}}}$, $^{\hat{\mathbf{y}}}$, and $^{\hat{\mathbf{z}}}$ are the orthonormal unit vectors in 3D space. The time-harmonic dependence $e^{\jmath\omega t}$, with $\omega$ denoting the angular frequency, is omitted for brevity. This current distribution acts as the source of the electromagnetic field, which propagates through space and interacts with the surrounding environment, including any scattering or receiving structures.
@@ III-A Preliminaries: The Green’s Operator and the Wavenumber Representation
In general, the space $\mathcal{S}_{\text{T}}$ represents a volume, although it may also correspond to a surface or a wire source. The current density ${\mathsf{J}}({\mathbf{r}})$ can be an impressed current in the case of an active source, such as a transmitting antenna, or an induced/equivalent current in the case of a scattering device. The electrical field ${\mathsf{E}}({\mathbf{r}})$ $\mathrm{V/m}$] generated by ${\mathsf{J}}({\mathbf{r}})$ can be computed under the free-space condition by considering the Helmholtz equation, whose solution is given by [35]
@@
${\mathsf{E}}({\mathbf{r}})=\int_{\mathcal{S}_{\text{T}}}\mathsf{G}\left({\mathbf{r}}-{\mathbf{s}}\right)\,{\mathsf{J}}({\mathbf{s}})\,\mathrm{d}{\mathbf{s}}=\left(\mathbb{G}\,{\mathsf{J}}\right)({\mathbf{r}})\,.$ (2)
@@
Equation (2) implicitly defines the Green’s tensor operator $\mathbb{G}$, with $\mathsf{G}\left({\mathbf{r}}\right)$ being the dyadic Green’s function defined as [35]
@@
$\mathsf{G}\left({\mathbf{r}}\right)=-\jmath\frac{\eta\,e^{-\jmath k_{0}\,|{\mathbf{r}}|}}{2\lambda|{\mathbf{r}}|}\cdot\left[\left({\mathbf{I}}-\hat{{\mathbf{r}}}\,\hat{{\mathbf{r}}}^{\mathrm{T}}\right)+\frac{\jmath\lambda}{2\pi|{\mathbf{r}}|}\left({\mathbf{I}}-3\hat{{\mathbf{r}}}\,\hat{{\mathbf{r}}}^{\mathrm{T}}\right)-\frac{\lambda^{2}}{(2\pi|{\mathbf{r}}|)^{2}}\left({\mathbf{I}}-3\hat{{\mathbf{r}}}\,\hat{{\mathbf{r}}}^{\mathrm{T}}\right)\right]\,$, (3)
@@
where $k_{0}=2\pi/\lambda$ is the wavenumber, $\lambda$ is the wavelength, $\hat{{\mathbf{r}}}={\mathbf{r}}/|{\mathbf{r}}|$, $\jmath$ denotes the imaginary unit, and $\eta$ is the free-space impedance. $\mathsf{G}\left({\mathbf{r}}\right)$ is the solution of the Helmholtz equation for an impulsive current density ${\mathsf{J}}({\mathbf{r}})={\mathsf{I}}\cdot\delta({\mathbf{r}})$, with $\mathsf{I}$} being the unit dyadic, [^1] The unit dyadic, also known as the identity dyadic or unit tensor, is a mathematical construct in vector and tensor analysis. It is the second-order tensor equivalent of the scalar number 1 in scalar algebra or the identity matrix in linear algebra. and $\delta({\mathbf{r}})$ is the delta Dirac pseudo-function. When $|{\mathbf{r}}|\gg\lambda$, only the first term in (3) contributes to the EM radiation. The electrical field in the current-free portion of the space must satisfy the homogeneous wave equation [35]
$\nabla^{2}{\mathsf{E}}({\mathbf{r}})+k_{0}^{2}\,{\mathsf{E}}({\mathbf{r}})=0\,$, (4)
@@
For what follows, it is convenient to introduce the representation of the fields in the wavenumber domain $^{\bm{\kappa}}=\kappa_{x}{\hat{\mathbf{x}}}+\kappa_{y}{\hat{\mathbf{y}}}+\kappa_{z}{\hat{\mathbf{z}}}$ through the 3D Fourier transform. Specifically, given a generic field $\mathsf{A}({\mathbf{r}})$, we define
$\widetilde{\mathsf{A}}({\bm{\kappa}})$ $=\mathcal{F}_{3}\left[\mathsf{A}({\mathbf{r}})\right]=\int_{{\mathbb{R}}^{3}}\mathsf{A}({\mathbf{r}})\,e^{-\jmath\,{\bm{\kappa}}\bm{\cdot}{\mathbf{r}}}\,\mathrm{d}{\mathbf{r}}\,$, (5)
[…]
@@
$\left(\kappa_{x}^{2}+\kappa_{y}^{2}+\kappa_{z}^{2}-k_{0}^{2}\right)\widetilde{\mathsf{E}}({\bm{\kappa}})=0\,$, (6)
@@
where we have considered that $\mathcal{F}_{3}\left[\nabla^{2}\mathsf{A}({\mathbf{r}})\right]=-|{\bm{\kappa}}|^{2}\,\widetilde{\mathsf{A}}({\bm{\kappa}})$. Equation (6) reveals that $\widetilde{\mathsf{E}}({\bm{\kappa}})=\mathcal{F}_{3}\left[{\mathsf{E}}({\mathbf{r}})\right]$ must vanish everywhere except on the wavenumber support
${\cal{E}}=\left\{(\kappa_{x},\kappa_{y},\kappa_{z})\in\mathbb{R}^{3}:\kappa_{x}^{2}+\kappa_{y}^{2}+\kappa_{z}^{2}=k_{0}^{2}\right\}\,.$ (7)
[…]
$\widetilde{\mathsf{E}}({\bm{\kappa}})=\widetilde{\mathsf{E}}_{0}({\bm{\kappa}})\,\delta\left(\kappa_{x}^{2}+\kappa_{y}^{2}+\kappa_{z}^{2}-k_{0}^{2}\right)\,$, (8)
@@
where $\widetilde{\mathsf{E}}_{0}({\bm{\kappa}})$ defines the angular dependence (direction) of the wave. When applied to the EM field ${\mathsf{E}}({\mathbf{r}})$ in the presence of current sources, the inverse Fourier representation expresses it in terms of mathematical plane waves [35, 15]
@@
${\mathsf{E}}({\mathbf{r}})$ $=\mathcal{F}^{-1}_{3}\left[\widetilde{\mathsf{E}}({\bm{\kappa}})\right]=\frac{1}{(2\pi)^{3}}\int_{{\mathbb{R}}^{3}}\widetilde{\mathsf{E}}({\bm{\kappa}})\,e^{\jmath\,{\bm{\kappa}}\bm{\cdot}{\mathbf{r}}}\,\text{d}{\bm{\kappa}}=\frac{1}{(2\pi)^{3}}\int_{{\mathbb{R}}^{3}}\widetilde{\mathsf{G}}({\bm{\kappa}})\,{\widetilde{\mathsf{J}}}({\bm{\kappa}})\,e^{\jmath\,{\bm{\kappa}}\bm{\cdot}{\mathbf{r}}}\,\text{d}{\bm{\kappa}}$
@@
$=\frac{\jmath\eta}{k_{0}(2\pi)^{3}}\int_{{\mathbb{R}}^{3}}\frac{{{\bm{\kappa}}}\,{\scriptstyle\mathsf{x}}\,{{{{\bm{\kappa}}}\,{\scriptstyle\mathsf{x}}\,{{\widetilde{\mathsf{J}}}({\bm{\kappa}})}}}}{|{\bm{\kappa}}|^{2}-k_{0}^{2}}\,e^{\jmath\,{\bm{\kappa}}\bm{\cdot}{\mathbf{r}}}\,\text{d}{\bm{\kappa}}\,$, (9)
where $\widetilde{\mathsf{G}}({\bm{\kappa}})=\mathcal{F}_{3}\left[\mathsf{G}\left({\mathbf{r}})\right)\right]$, ${\widetilde{\mathsf{J}}}({\bm{\kappa}})=\mathcal{F}_{3}\left[{\mathsf{J}}({\mathbf{r}})\right]$, and $\scriptstyle\mathsf{x}$}\,{} denotes the cross product. In fact, for a fixed wavenumber $\bm{\kappa}$}, $\widetilde{\mathsf{E}}({\bm{\kappa}})\,e^{\jmath\,{\bm{\kappa}}\bm{\cdot}{\mathbf{r}}}$ in (9) represents a plane wave with direction ${\bm{\kappa}}/k_{0}$.
@@
It is interesting to emphasize that the generic plane wave $\widetilde{\mathsf{E}}({\bm{\kappa}})\,e^{\jmath\,{\bm{\kappa}}\bm{\cdot}{\mathbf{r}}}=\widetilde{\mathsf{E}}({\bm{\kappa}})\,e^{\jmath\,\kappa_{x}\,r_{x}}\,e^{\jmath\,\kappa_{y}\,r_{y}}\,e^{\jmath\,\kappa_{z}\,r_{z}}$ in (9) radiates $\forall{\mathbf{r}}$, i.e., in all the 3D space, only when $\kappa_{x}$, $\kappa_{y}$, and $\kappa_{z}$ are all real (visible region). Vice versa, if at least one of them is complex, [^3] For example, this might happen if $\kappa_{x}^{2}+\kappa_{y}^{2}>k_{0}^{2}$ so that to satisfy (7) it should be $\kappa_{z}=\pm\jmath\sqrt{\kappa_{x}^{2}+\kappa_{y}^{2}-k_{0}^{2}}\in\mathbb{C}$. let us say $\kappa_{z}$, it is $e^{\jmath\,\kappa_{z}\,r_{z}}=e^{-|\kappa_{z}\,r_{z}|}\rightarrow 0$ for $|r_{z}|$ larger than a few wavelengths and the wave becomes evanescent, i.e., it does not radiate (invisible region). In summary, only the spectral components of $\widetilde{\mathsf{E}}({\bm{\kappa}})$ for $\bm{\kappa}$$\in{\cal{E}}$ represent EM waves, i.e., are physically consistent, and only the subset having real components radiates. For each wavenumber, i.e., plane wave, because of the ${{\bm{\kappa}}}\,{\scriptstyle\mathsf{x}}\,{{{{\bm{\kappa}}}\,{\scriptstyle\mathsf{x}}\,{{\widetilde{\mathsf{J}}}({\bm{\kappa}})}}}$ term in (9), $\widetilde{\mathsf{E}}({\bm{\kappa}})$ is subject to the relation $\widetilde{\mathsf{E}}({\bm{\kappa}})\bm{\cdot}{\bm{\kappa}}=0$, limiting to two the DoF achievable with different polarizations [35]. Incidentally, for locations $\mathbf{r}$} in the far-field region of the source, (9) can be approximated as [35]
${\mathsf{E}}({\mathbf{r}})$ $\simeq\jmath k_{0}\eta\frac{e^{-\jmath\,k_{0}\,|{\mathbf{r}}|}}{4\pi|{\mathbf{r}}|}\,{\hat{\mathbf{r}}}\,{\scriptstyle\mathsf{x}}\,{\,{{\hat{\mathbf{r}}}\,{\scriptstyle\mathsf{x}}\,{{\widetilde{\mathsf{J}}}(k_{0}\,\hat{\mathbf{r}})}}}\,$, (10)
@@
so that the EM field becomes directly related to the Fourier transform of the source evaluated at direction $k_{0}\,\hat{\mathbf{r}}\in{\cal{E}}$.